A motel rents rents double rooms at $75 per night and single rooms at $64 per night. one night, the income from renting double and single rooms was $2,244. if the number of single rooms rented was three less than two times the number of rooms rented, how many rooms of each kind were rented?
Answer by mananth(16949)

I think there is a typo
it should be 3 less than twice the number of double rooms

single rooms rented be x
double rooms rented =y
x= 2y-3
x=2y-3
x=2y-3............(1)

cost
75y+64x=2244..............(2)
substitute x in (2)
75y+64(2y-3)=2244
75y+128y-192=2244
203y=2244+192
203y=2436
y=12 the number of double rooms
x=2y-3
plug y
x=2*12-3
x=21 single rooms
m.ananth@hotmail.ca